may you please hepe me : In the euromillion lottery,a player picks five numbers from a choice of 1 to 50 and two lucky star number from 1 to 11. the jackpot is awarded to any player who matches the five numbers and two lucky stars. What is the probability of winning the jackpot? (I've used the binomial formula to do this question and ive got an answer of 2.79x10^10, im not sure how to layout my working out.) thank you so much.
Answer by stanbon(75887)

may you please hepe me : In the euromillion lottery,a player picks five numbers from a choice of 1 to 50 and two lucky star number from 1 to 11. the jackpot is awarded to any player who matches the five numbers and two lucky stars. What is the probability of winning the jackpot?
-------
Prob of picking the 5 winning numbers = 1/50C5
-------
Prob of picking the 2 winning star numbers = 1/11C2
-------
P(winning the lottery) = (1/50C5)(1/11C2) = (1/2118760)*(1/55)
= 8.58x10^-9
-----
= 0.0000000858
----
That's 858 out of 10 billion.
----
Save your money.
====================
Cheers,
Stan H.
============